College tuition: a simple random sample of 35 colleges and universities in the united states has a mean tuition of $19,400 with a standard deviation of $10,700. Construct a 95% confidence interval for the mean tuition for all colleges and universities in the United States. round the answers to the nearest whole number.

A 95% confidence interval for the mean tuition for all colleges and universities is __________________< mean < _____________________
Answer #1

Solution :

Given that,

\bar x = $19400

s = $10700

n = 35

Degrees of freedom = df = n - 1 = 35 - 1 = 34

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,34 = 2.032

Margin of error = E = t\alpha/2 * (s /\sqrtn) = 2.032 * (10700/ \sqrt35) = 3675

A 95% confidence interval for the mean tuition for all colleges and universities is

\bar x - E < \mu < \bar x + E

19400 - 3675 < \mu < 19400 + 3675

$15725 < \mu < $23075